# Aligning Large Language Models for Controllable Recommendations

### v1

@@ 3.2.2 Category Control
$\mathrm{Ours_{v2}}$ 0.2620 0.1814 $81.38(8.05)$ 0.1099 0.0832 $\underline{19.35}(39.80)$
$\mathrm{Ours_{v3}}$ $0.2383$ $0.1609$ $\underline{89.06}(7.89)$ $0.1017$ $0.0772$ $15.80(39.04)$
$\mathrm{Ours_{full}}$ $\underline{0.2336}$ $\underline{0.1574}$ $\textbf{93.52}(7.81)$ $0.0999$ $0.0756$ $\textbf{11.49}(38.41)$
@@ 3.2.3 Category Proportion Control
Results are reported in Table 4. In this case, the primary metric is CPA. Overall, $Ours_{full}$ achieves the best performance, outperforming all other baselines and its own variants.
@@
$\mathrm{Ours_{v2}}$ 0.1027 $\underline{0.0688}$ $29.18(14.93)$ 0.2034 0.1417 $\underline{58.71}(4.61)$ $\underline{0.2055}$ $\underline{0.1433}$ $45.89(9.96)$
$\mathrm{Ours_{v3}}$ $0.0943$ $0.0594$ $\underline{33.74}(15.03)$ $0.1948$ $0.1319$ $65.78(4.58)$ $0.1994$ $0.1340$ $48.19(9.21)$
@@ 3.3 Formatting and General Evaluation
Table 5 presents the comprehensive outcomes. Notably, nearly all models, with the exception of PALR on the Movie dataset, excel in the CorrectCount metric. For other formatting metrics such as RepeatItem, NonExist, and InHistory, a progressive enhancement is evident from $Ours_{v1}$ through $Ours_{v3}$, culminating in $Ours_{full}$. In Table 5, the primary focus is on Formatting Quality, while Precision metrics serve as supplementary references. The Generalization metrics indicate the extent of catastrophic forgetting. When compared to the fine-tuned baseline PALR, $Ours_{full}$ shows the least performance deterioration from its underlying Llama2-7b architecture. Owing to the 512-token input limitation of InstructRec, its performance could not be assessed on the Generalization tasks.
@@ 3.4 Case Study
For clarity, we present two illustrative cases. The first is about the $I_{2}^{CP\approx 50\%}$ instruction on the Movie dataset. A user requests that roughly 50% of the top-5 recommendations feature “Art House & International, By Original Language, Chinese” characteristics. Our method successfully generates a list where 3 out of 5 items possess these attributes, while three baseline methods fail to meet this criterion. Llama-2-7b simply repeats items mentioned in the user history.
@@ 4 Related Work
Initially, adaptation of LLMs to recommendation scenarios relied heavily on techniques like prompt engineering Gao et al. (2023); Sun et al. (2023) and contextual learning Dai et al. (2023); Liu et al. (2023). However, these LLMs often underperform in comparison to traditional recommendation models trained on task-specific data, prompting efforts to fine-tune LLMs for better alignment with recommendation tasks. P5 Geng et al. (2022) introduced a unified framework that integrates five recommendation tasks through fine-tuning on the FLAN-T5 model Raffel et al. (2020). Subsequently, InstructRec Zhang et al. (2023) tailored FLAN-T5 for various downstream recommendation tasks using instruction tuning. TALLRec Bao et al. (2023) fine-tunes the LLaMA model for binary recommendation tasks with limited training samples, while PALR Chen (2023) employs two types of instructions for instruction tuning to facilitate list-level recommendation generation.
@@ 5 Conclusion
In conclusion, we introduce a novel alignment methodology for LLMs in recommender systems, which significantly enhances their ability to follow user instructions and reduces formatting errors. By incorporating a supervised learning stage with a novel suite of control tasks and a reinforcement learning stage tailored with custom reward signals, our approach effectively bridges the gap between LLM capabilities and the nuanced requirements of recommendation tasks. Experiments confirm that our method not only outperforms existing LLM-based recommender systems but also demonstrates promising capabilities towards controllable and accurate recommendation experiences.
@@ 6 Limitation
This paper’s primary constraints are as follows: (1) The emphasis is placed on improving the LLM’s ability to follow recommendation-related instructions. This focus may inadvertently compromise the LLM’s general capabilities, as revealed in Table 5. (2) In real-world scenarios, users often have diverse control intentions, including intricate blends of various instructions. As foundational research, this paper addresses only the most critical elements, specifically category control and formatting control.
@@ A.1 Dataset detail
We present detailed information about the instruction set we constructed in Table 7, including the quantity of each type of instruction.
@@ 31      $Scores^{ctl}_{i}=s$
A.2 Prompts

### v2

@@ 3.2.2 Category Control
$\mathrm{Ours_{v2}}$ 0.2620 0.1814 $81.38(8.05)$ 0.1099 0.0832 $19.35(39.80)$
$\mathrm{Ours_{v3}}$ $\underline{0.2383}$ $\underline{0.1609}$ $\underline{89.06}(7.89)$ $0.1017$ $0.0772$ ${15.80}(39.04)$
$\mathrm{Ours_{full}}$ $0.2336$ $0.1574$ $\textbf{93.52}(7.81)$ $0.0999$ $0.0756$ $\underline{11.49}(38.41)$
@@ 3.2.3 Category Proportion Control
Results are reported in Table 4. In this case, the primary metric is CPA. Overall, Oursfull achieves the best performance, outperforming all other baselines and its own variants.
@@
$\mathrm{Ours_{v2}}$ 0.1027 $\underline{0.0688}$ $29.18(14.93)$ 0.2034 0.1417 $58.71(4.61)$ $\underline{0.2055}$ $\underline{0.1433}$ $45.89(9.96)$
$\mathrm{Ours_{v3}}$ $0.0943$ $0.0594$ $\underline{33.74}(15.03)$ $0.1948$ $0.1319$ $\underline{65.78}(4.58)$ $0.1994$ $0.1340$ $48.19(9.21)$
@@ 3.3 Formatting and General Evaluation
Table 5 presents the comprehensive outcomes. Notably, nearly all models, with the exception of PALR on the Movie dataset, excel in the CorrectCount metric. For other formatting metrics such as RepeatItem, NonExist, and InHistory, a progressive enhancement is evident from Oursv1 through Oursv3, culminating in Oursfull. In Table 5, the primary focus is on Formatting Quality, while Precision metrics serve as supplementary references. The Generalization metrics indicate the extent of catastrophic forgetting. When compared to the fine-tuned baseline PALR, Oursfull shows the least performance deterioration from its underlying Llama2-7b architecture. Owing to the 512-token input limitation of InstructRec, its performance could not be assessed on the Generalization tasks.
@@ 3.4 Case Study
For clarity, we present two illustrative cases in Table 6. The first is about the $I_{2}^{CP\approx 50\%}$ instruction on the Movie dataset. A user requests that roughly 50% of the top-5 recommendations feature “Art House & International, By Original Language, Chinese” characteristics. Our method successfully generates a list where 3 out of 5 items possess these attributes, while three baseline methods fail to meet this criterion. Llama-2-7b simply repeats items mentioned in the user history.
@@
$\bm{Control}$ $I_{1}^{+C_{1}}\ \&\ I_{1}^{-C_{2}}$ $I_{2}^{C_{1}P\leq 20\%}\ \&\ I_{1}^{-C_{2}}$
Dataset Model $\mathrm{HR@10}$ $\mathrm{TC_{1\neg 2}P@10(\%)}\uparrow$ $\mathrm{HR@10}$ $C_{1}PA\uparrow$ $\mathrm{TC_{2}P@10(\%)}\downarrow$
Movie $\mathrm{GPT-3.5}$ $0.0176$ $7.15(2.20)$ $0.0132$ $9.84(6.37)$ $\textbf{12.51}(19.45)$
[…]
$\mathrm{InstructRec}$ $0.0944$ $27.23(7.14)$ $0.0333$ $12.83(9.47)$ $51.81(31.91)$
$\mathrm{PALR}$ $0.0762$ $7.99(7.60)$ $0.0692$ $33.21(6.79)$ $38.08(53.46)$
$\mathrm{Ours_{v1}}$ $0.1002$ $8.21(8.05)$ $0.0965$ $28.40(11.94)$ $28.69(31.27)$
$\mathrm{Ours_{v2}}$ $0.2031$ $34.63(8.04)$ $\underline{0.1012}$ $29.13(13.21)$ $25.28(30.40)$
$\mathrm{Ours_{v3}}$ $\underline{0.2033}$ $\underline{44.16}(7.88)$ $0.0927$ $\underline{36.25}(15.07)$ $22.01(28.88)$
$\mathrm{Ours_{full}}$ 0.2064 $\textbf{49.61}(7.89)$ 0.1018 $\textbf{45.14}(15.95)$ $\underline{16.65}(28.50)$
Steam $\mathrm{GPT-3.5}$ $0.0194$ $34.15(21.85)$ $0.0183$ $15.38(4.21)$ $\underline{22.13}(34.18)$
[…]
$\mathrm{InstructRec}$ $0.0432$ $27.94(18.63)$ $0.0133$ $13.84(20.37)$ $38.23(34.63)$
$\mathrm{PALR}$ $0.0353$ $10.52(11.09)$ $0.0294$ $7.04(2.53)$ $41.31(47.31)$
$\mathrm{Ours_{v1}}$ $0.0859$ $25.07(25.31)$ $0.0842$ $21.58(12.25)$ $36.18(41.81)$
$\mathrm{Ours_{v2}}$ $0.2620$ $42.39(26.95)$ $\underline{0.1187}$ $54.06(14.58)$ $29.63(39.34)$
$\mathrm{Ours_{v3}}$ $\underline{0.2729}$ $\underline{48.27}(28.83)$ $0.1183$ $\underline{75.56}(17.74)$ $22.38(35.74)$
$\mathrm{Ours_{full}}$ 0.2807 $\textbf{56.00}(28.99)$ 0.1204 $\textbf{77.26}(17.81)$ $\textbf{18.82}(35.56)$
Table 7: Results of two combinatorial control instructions: (1) $I_{1}^{+C_{1}}\ \&\ I_{1}^{-C_{2}}$ ; (2) $I_{2}^{C_{1}P\leq 20\%}\ \&\ I_{1}^{-C_{2}}$. The best result is highlighted in boldface and the runner-up is denoted with underline. Values in parentheses show outcomes of the corresponding model without control signals.
3.5 Extend to combinatorial Controls
Considering LLMs’ strong generalization capabilities, we further examine the performance of LLMs on complex, combinatorial instructions not encountered during training, as shown in Table 7. Here, $TC_{1\neg 2}P@10(\%)$ means the percentage of items belong to $C_{1}$ but does not belong to $C_{2}$. The results confirm our expectations that while our methods surpass baseline models, they achieve lower compliance with user instructions, as measured by TCP and CPA, compared to simpler single-control instructions. This motivates us to include some complex instructions in the alignment process in future work.
@@ 4 Related Work
Initially, adaptation of LLMs to recommendation scenarios relied heavily on techniques like prompt engineering Gao et al. (2023); Sun et al. (2023) and contextual learning Dai et al. (2023); Liu et al. (2023). However, these LLMs often underperform compared to traditional recommendation models trained on task-specific data, prompting the necessity to fine-tune LLMs for better alignment with recommendation tasks. P5 Geng et al. (2022) introduced a unified framework that integrates 5 recommendation tasks through fine-tuning on the FLAN-T5 model Raffel et al. (2020). Subsequently, InstructRec Zhang et al. (2023) tailored FLAN-T5 for various downstream recommendation tasks using instruction tuning. TALLRec Bao et al. (2023) fine-tunes LLaMA for recommendations with very few training samples, but it focuses on the binary classification task. PALR Chen (2023) employs two types of instructions for instruction tuning to facilitate list-level recommendation generation.
@@ 5 Conclusion
In conclusion, our paper presents a new approach to tailor LLMs for interactive recommender systems, combining a supervised learning phase with innovative control tasks and a reinforcement learning stage with specialized reward signals. Our method successfully meets the detailed demands of recommendation contexts, enhancing LLMs’ performance. Experimental results show our approach exceeds current LLM-based systems in precision, controllability, and presentation, offering a significant step towards refined and reliable recommendation services.
@@ 6 Limitation
This paper’s primary constraints are as follows: (1) The emphasis is placed on improving the LLM’s ability to follow recommendation-related instructions. This focus may inadvertently compromise the LLM’s broader intellectual capabilities, as revealed in Table 5. How to further reduce catastrophic forgetting remains a big challenge. (2) In real-world scenarios, users often have diverse control intentions, including intricate blends of various instructions (as discussed in Section 3.5) and new instructions beyond category control. As foundational research, this paper addresses only the most critical elements, specifically category control and formatting control. More diverse and complicated instructions are yet to be explored.
7 Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant Nos. 62276171 and 62072311), Guangdong Basic and Applied Basic Research Foundation (Grant Nos. 2024A1515011938, 2020B1515120028), Guangdong Provincial Key Laboratory of Popular High Performance Computers (Grant No. 2017B030314073), Guangdong Province Engineering Center of China-made High Performance Data Computing System, Shenzhen Fundamental Research-General Project (Grant Nos. 20220811155803001, 20210324094402008). Hao Liao is the corresponding author.
@@ A.1 Dataset detail
We present detailed information about the instruction set we construct in Table 8, including the quantity of each type of instruction.
@@ 31      $Scores^{ctl}_{i}=s$
A.2 Computational Cost
Finetuning LLMs for recommendations is a little bit more expensive than training traditional recommender models like SASRec, but the resource and cost are not a big barrier even for small organizations or researchers. Here we take the Movie dataset for illustration. The dataset statistics can be found in Table 1 and Table 7.
For the SL stage, we use 4 A100 GPUs (40GB GPU memory) for training. The batch size on each GPU is 1 and the gradient accumulation step of 16. The maximum sequence length is set to 1024. Under this setting, each epoch costs about 70 minutes. Usually, models can fully converge in 30 epochs. So, the total training time for SFT is about 35 hours.
For the RL stage, we use 2 A100 GPUs (40GB GPU memory) for training. The batch size on each GPU is 1 and the gradient accumulation step of 2. For each batch, instruction sampling time + 2 candidate response generation time + model training time counts about 40 seconds, and the maximum training steps are set to 3k. Thus, the total RL training time is $3000*40/60/60=33$ hours.
Once the training process is finished, we use vllm [^5] https://github.com/vllm-project/vllm for inference and serving. On a single A100 GPU (40GB memory), when responding to the top-10 recommendation request, our implementation can process 20 requests per second (on average each request involves newly generated 112 tokens).
A.3 Prompts